### v1

@@ 1 Introduction
Machine learning (ML) has recently emerged as a powerful alternative in materials research and is gaining traction, with several disciplines employing these techniques to expedite the discovery of new materials [16, 17, 18, 19, 20, 21, 13, 22, 23]. Data-driven algorithms can analyze vast datasets of catalyst properties and performance, identifying complex relationships that may be beyond the reach of traditional descriptors [24, 25]. In the realm of heterogeneous catalysis, the ML methods can primarily be divided into two categories: mapping catalyst activity using new approximate descriptors [18, 20, 22] and prediction of adsorption energies using machine-learned force fields (MLFF) [17, 21]. These MLFFs, trained on extensive datasets of DFT calculations, offer a significant speedup (a factor of 104 or more) compared to DFT calculations while maintaining accuracy [26, 27]. Modern Sabatier principle-based approaches utilize MLFFs or specialized models to find (global) minimum adsorption energies across multiple material facets [21, 13]. Although several studies emphasize the importance of various catalyst facets [9, 28], these approaches predominantly utilize data from individual facets for material characterization. Consequently, the challenge persists: how can we effectively predict catalytic performance without limiting our scope to specific material families or facet orientations?
The absence of an adequate descriptor for the activity of complex materials motivates our exploration of methods to better represent contemporary industrial catalysts. These catalysts, composed of nanostructures with diverse surface facets and adsorption sites, present significant challenges in understanding their performance. This work seeks to address these challenges by focusing on three critical objectives: (1) developing a novel descriptor that captures the structural and energetic complexity of catalysts, (2) establishing an efficient workflow for large-scale computational screening, and (3) devising a robust framework to identify promising candidates from the resulting data.
@@
Secondly, we aim to establish a high-throughput and ML-enhanced workflow to accelerate the screening of catalytic materials using our newly formulated descriptor. Traditional density functional theory (DFT) approaches are computationally prohibitive for large-scale studies. To overcome this limitation, we leverage machine-learned force fields (MLFFs) from the Open Catalyst Project (OCP) [31, 32], enabling the rapid and accurate computation of adsorption energies. Our workflow generates an extensive dataset of AEDs, capturing over 877,000 adsorption energies for nearly 160 materials relevant to the CO2 to methanol conversion reaction. To target the enhanced reliability of our prediction and the effective use of ML models in catalyst discovery, we design a robust validation protocol.
@@ 2 Results and Discussion
To identify the most crucial adsorbates for AEDs calculations, we perused the existing literature. An experimental investigation by Amman et al. [28] highlighted the presence of surface-bound radicals such as *H, *OH, *OCHO, and *OCH3 as essential reaction intermediates in the thermocatalytic reduction of CO2 to methanol. Based on these findings, we selected them for our AEDs calculations. With the help of fairchem repository tools by OCP [37], we created surfaces with their Miller index $\in\{-2,-1,...,2\}$ and calculated their total energy using OCP MLFF. If we encountered multiple cuts for the same facet, we selected the one with lowest energy for further calculations. Then we engineered surface-adsorbate configurations for the most stable surface terminations across all facets within our defined Miller index range for the materials, as described in section 3.2, and optimized these configurations using the OCP MLFF. During this process, we discovered that six materials exhibited so large surface-adsorbate supercells, that their calculations were infeasible on available GPU resources, even with the effective OCP MLFF. Consequently, they were excluded from our study.
@@
Figure 3: AEDs for selected subset of materials: (a) AEDs for 26 materials with the amplitude of the distribution encoded in the intensity of the corresponding color; detailed AEDs for (b) K, (c) Y3In5, (d) Ni, (e) CuZn, (f) NiZn, (g) ZnRh.
@@
Although the ideal AEDs for the four adsorbates remain unknown, it is feasible to approximate their reactivity using AEDs based on their resemblance to previously identified, efficacious catalytic materials. In this context, our AEDs can be conceptualized as four-dimensional probability distributions. To quantify similarities across AEDs of different materials, we employ the Wasserstein distance as the metric [33]. By computing Wasserstain distances for all possible material combinations, we construct a distance matrix. To interpret the distance matrix, we apply hierarchical agglomerative clustering with Ward’s linkage [40], which facilitates the identification of materials with similar AEDs. The outcomes of this clustering analysis are depicted in Fig. 4.
@@
Clusters 8 through 10 aggregate into a larger cluster with relatively homogeneous AEDs, encompassing materials such as Cu, a notably active component within known Cu/ZnO/Al2O3 catalysts [28, 34]. The clusters also contain non-Cu materials such as Zn-Pd, Pd-In, Pt-In, and Ni-Zn in different compositions that have been reported as catalytic converters of CO2 to methanol [34, 41]. Several bimetallic alloys are grouped with these materials that have not been tested for CO2 to methanol conversion, up to our knowledge: Ga-Ag, In-Ag, K-In, Zn-Rh, and Zn-Pt, which also occur in various compositions. While most of these materials have either shown good catalytic performance or some have not been tested, potassium (K) (the lone non-numbered cluster), as a pure metal, is likely to undergo rapid oxidation under reaction conditions. Therefore, we anticipate that this particularly large cluster, consisting of clusters 8-10, is likely too diverse to pinpoint only catalytically active materials.
Upon closer inspection, ZnRh and ZnPt3 stand out as new candidates. They are part of cluster 10, which also includes Ga2Cu, NiZn, InPt3, Ni and mainly CuZn, but have not been tested for CO2 to methanol conversion. It is evident that Zn incorporation into Cu enhances the activity of Cu-based catalysts, predominantly due to the formation of a Cu-Zn alloy [42, 43, 28]. NiZn has also been identified as an effective CO2 catalyst [41]. Catalysts such as Cu/Ga2O3 and Pt/In2O3, known for their high methanol yield, likely include Ga2Cu and InPt3 alloys, respectively. Finally, Ni is often part of catalysts for CO2 transformation to methane [44, 45]. The strong catalytic activity of Ga2Cu, NiZn, InPt3, Ni and mainly CuZn in this cluster suggest that also ZnRh and ZnPt3 should have a high activity.
@@
Both catalyst descriptors are tailored for an extensive search for catalytically active material candidates, yet they omit the effect of support, additives, and preparation procedures that can change the size of the active material nanoparticles and the area of different facets. Moreover, our AED descriptor does not take the facet area into account, and is therefore insensitive to morphology changes of the catalysts under reaction conditions.
To finalize the analysis of our results, the similarity of the SAAED and Wasserstein distances of ZnRh and ZnPt3 to good catalysts in the literature suggests that they could be good catalyst candidates. As Cu-based catalysts are known for their vulnerability to degradation [5], it is therefore reasonable to pre-examine these materials also in terms of stability. Given the harsh reaction conditions, mainly temperatures around 800 K [34], the melting temperature of the catalyst is directly related to the stability of the catalyst. The melting temperature of both ZnRh and ZnPt3 is higher than that of pure copper or CuZn [35], suggesting that our new candidates could also be more durable.
@@ 3.3 Generation and Relaxation of Adsorbate-Surface Configurations
The selected surfaces were used to generate adsorbate-surface configurations using neutral fragments of *H, *OH, *OCHO and *OCH3 by means of the fairchem input generation workflow [37]. The systems were relaxed using an adsorption energy-based OCP equiformer_V2 MLFF to obtain the relaxed geometries. The corresponding adsorption energy $E_{ads}$ follows the OCP convention $E_{ads}=E_{system}-E_{surface}-E_{adsorbate}$, where $E_{adsorbate}$ is composed of atomic energies, defined in the work of Chanussot et. al. [31]. Here, $E_{adsorbate}$ is –3.477 eV for *H, –10.552 eV for *OH, –24.917 eV for *OCH3 and –25.161 eV for *OCHO. The force convergence criterion for the adsorbate relaxations was set to 0.03 eV/Å.
The equiformer_V2 relaxations were performed using NVIDIA Ampere A100 GPUs, completing each relaxation in just a few seconds, whereas corresponding DFT calculations would require several hours on two AMD Rome 7H12 CPUs with 64 cores each. Relaxed configurations of the same material and facet were compared to identify redundancies. If two or more configurations exhibited identical adsorbate geometries (and thus energies) within a tolerance of 0.1 Å in each spatial direction, only one configuration was retained for further consideration. Nonunique configurations and their corresponding adsorption energies were excluded. With this we ensured that the remaining data represents meaningful adsorption positions and thus capture the material’s true energetic landscape.

### v2

@@ 1 Introduction
Machine learning (ML) has recently emerged as a powerful alternative in materials research and is gaining traction, with several disciplines employing these techniques to expedite the discovery of new materials [16, 17, 18, 19, 20, 21, 13, 22, 23]. Data-driven algorithms can analyze vast datasets of catalyst properties and performance, identifying complex relationships that may be beyond the reach of traditional descriptors [24, 25]. In the realm of heterogeneous catalysis, the ML methods can primarily be divided into two categories: mapping catalyst activity using new approximate descriptors [18, 20, 22, 26] and prediction of adsorption energies using machine-learned force fields (MLFF) [17, 21].
Both of these approaches have significantly boosted effectiveness in heterogeneous catalyst research [27]. Descriptor-based approaches such as SISSO [18] and the latest generation of MLFFs based on graph neural networks [21], offer remarkable precision. However, these methods require training, which results in a computational cost comparable to direct DFT calculations.
In contrast, pre-trained MLFFs, that make use of extensive DFT datasets, offer explicit relaxation of adsorbates on catalyst surfaces and a significant speed-up (a factor of 104 or more) compared to DFT calculations while maintaining quantum mechanical accuracy [28, 29] making them suitable for high-throughput catalyst discovery workflows. Modern Sabatier principle-based approaches also utilize MLFFs or specialized models to find (global) minimum adsorption energies across multiple material facets [21, 13] providing a scalable and transferable framework for screening a broad range of catalysts. Although several studies emphasize the importance of various catalyst facets [9, 30], these approaches predominantly utilize data from individual facets for material characterization. Consequently, the challenge persists: how can we effectively predict catalytic performance without limiting our scope to specific material families or facet orientations?
The absence of an adequate descriptor for the activity of complex materials motivates our exploration of methods to better represent contemporary industrial catalysts. These catalysts, composed of nanostructures with diverse surface facets and adsorption sites, present significant challenges to understanding their performance. This work seeks to address these challenges by focusing on three critical objectives: (1) developing a novel descriptor that captures the structural and energetic complexity of catalysts, (2) establishing an efficient workflow for large-scale computational screening, and (3) devising a robust framework to identify promising candidates from the resulting data.
@@
Secondly, we aim to establish a high-throughput and ML-enhanced workflow to accelerate the screening of catalytic materials using our newly formulated descriptor. Traditional density functional theory (DFT) approaches are computationally prohibitive for large-scale studies. To overcome this limitation, we leverage MLFFs from the Open Catalyst Project (OCP) [33, 34], enabling the rapid and accurate computation of adsorption energies. Our workflow generates an extensive dataset of AEDs, capturing over 877,000 adsorption energies for nearly 160 materials relevant to the CO2 to methanol conversion reaction. To target the enhanced reliability of our prediction and the effective use of ML models in catalyst discovery, we design a robust validation protocol.
@@ 2 Results and Discussion
To identify the most crucial adsorbates for AEDs calculations, we perused the existing literature. An experimental investigation by Amman et al. [30] highlighted the presence of surface-bound radicals such as *H (hydrogen atom), *OH (hydroxy group), *OCHO (formate), and *OCH3 (methoxy) as essential reaction intermediates in the thermocatalytic reduction of CO2 to methanol. Based on these findings, we selected them for our AEDs calculations. Please note that the notation for formate can vary across the literature, e.g., *HCOO [30, 39] and HCOO* [40]. With the help of fairchem repository tools by OCP [41], we created surfaces with their Miller index $\in\{-2,-1,...,2\}$ and calculated their total energy using OCP MLFF. If we encountered multiple cuts for the same facet, we selected the one with lowest energy for further calculations. Then we engineered surface-adsorbate configurations for the most stable surface terminations across all facets within our defined Miller index range for the materials, as described in section 3.2, and optimized these configurations using the OCP MLFF. During this process, we discovered that six materials exhibited so large surface-adsorbate supercells, that their calculations were infeasible on available GPU resources, even with the effective OCP MLFF. Consequently, they were excluded from our study.
@@
Figure 3: AEDs for selected subset of materials: (a) AEDs for 26 material with the amplitude of the distribution encoded in the intensity of the corresponding color; detailed AEDs for (b) K, (c) Y3In5, (d) Ni, (e) CuZn, (f) NiZn, (g) ZnRh. The numbers on the y-axis of panel (a) indicate the cluster numbers defined in Fig. 4.
@@
Although the ideal AEDs for the four adsorbates remain unknown, it is feasible to approximate their reactivity using AEDs based on their resemblance to previously identified, efficacious catalytic materials. In this context, our AEDs can be conceptualized as four-dimensional probability distributions. To quantify similarities across AEDs of different materials, we employ the Wasserstein distance as the metric [35]. By computing Wasserstein distances for all possible material combinations, we construct a distance matrix. To interpret the distance matrix, we apply hierarchical agglomerative clustering with Ward’s linkage [44], which facilitates the identification of materials with similar AEDs. The outcomes of this clustering analysis are depicted in Fig. 4.
@@
Clusters 8 through 10 aggregate into a larger cluster, from hereon denoted macro-cluster, with relatively homogeneous AEDs. It encompasses materials such as Cu, a notably active component within known Cu/ZnO/Al2O3 catalysts [30, 36]. The clusters also contain non-Cu materials such as Zn-Pd, Pd-In, Pt-In, and Ni-Zn in different compositions that have been reported as catalytic converters of CO2 to methanol [36, 45]. Also different compositions of the bimetallic alloys Ga-Ag, In-Ag, K-In, Zn-Rh, and Zn-Pt that, to our knowledge, have not been tested for CO2 to methanol conversion are grouped with these materials. While most of the materials in the macro-cluster have either shown good catalytic performance or some have not been tested, potassium (K) (the lone non-numbered cluster), as a pure metal, is likely to undergo rapid oxidation under reaction conditions. Therefore, we anticipate that the macro-cluster, consisting of clusters 8-10, is likely too diverse to pinpoint only catalytically active materials.
Upon closer inspection, ZnRh and ZnPt3 stand out as new candidates. They are part of cluster 10, which also includes Ga2Cu, NiZn, InPt3, Ni and mainly CuZn, but have not been tested for CO2 to methanol conversion. While the exact material composition of the Cu and ZnO-based catalysts during the proceeding reaction is still debated [46, 47, 48], some studies suggest that the formation of a Cu-Zn alloy enhances the activity. This Cu-Zn alloy, which is part of cluster 10, is believed to contribute to increased activity [47, 49, 30]. Similarly, NiZn has also been identified as an effective CO2 catalyst [45]. Catalysts such as Cu/Ga [50] and Pt/In2O3 [51], known for their high methanol yield, may include Ga2Cu and InPt3 alloys, respectively. Finally, Ni is often part of catalysts for CO2 transformation to methane [52, 53]. The strong catalytic activity of Ga2Cu, NiZn, InPt3, Ni and mainly CuZn in this cluster suggest that also ZnRh and ZnPt3 should have a high activity.
To conclude this section, we reiterate that our approach groups catalyst materials according to their computed AED similarity. To ascribe meaning to certain similarities we observe, we currently rely on experimental reports of the catalytically active materials. Our proposals for interesting candidates are based on the assumption that AED similarity with a known good catalyst is a meaningful indicator for promising catalytic activity. Since the catalyst composition and microstructure are often not reported or not known, “active sites” or details of the catalytic mechanisms also remain opaque [30, 54, 48, 46]. In this context, our AED descriptor remains an attempt to find proxies for complex processes. It goes beyond the common practice of focusing on single adsorption energies in “active sites”, but could certainly be extended in future work and in collaboration with more detailed experimental investigations.
@@
Both catalyst descriptors are tailored to perform extensive searches for catalytically active materials candidates. They do, however, not include effects of the support, additives, preparation procedures and operando states that could change the morphology of the catalyst (e.g., nanoparticle sizes or areas of different facets). Moreover, our AED descriptor does not take the facet area into account, and is therefore insensitive to morphology changes of the catalysts under reaction conditions. In principle, a Wulff construction could introduce better facet information. However, it also cannot account for support effects, additives or preparation conditions. Incorporating general descriptors (e.g., Magpie) for support materials could provide additional information to decide which active material-support combination should be investigated further in experimental testing
Our workflow clusters materials with high CO2 conversion efficiency, but the materials can vary in their selectivity towards methanol or methane [50, 36, 51, 45, 55]. As the reaction conditions, preparation procedures, or the interaction with support materials seem to affect the selectivity [45, 55], our proposed catalyst candidates should thus be tested under various conditions to investigate their optimal selectivity towards methanol.
To finalize the analysis of our results, the similarity of the SAAED and Wasserstein distances of ZnRh and ZnPt3 to good catalysts in the literature suggests that they could be potential catalyst candidates. As Cu-based catalysts are known for their vulnerability to degradation [5], it is therefore reasonable to pre-examine these materials also in terms of stability. Given the harsh reaction conditions, mainly temperatures around 800 K [36], the melting temperature of the catalyst is directly related to the stability of the catalyst. The melting temperature of both ZnRh and ZnPt3 is higher than that of pure copper or CuZn [37], suggesting that our new candidates could also be more durable.
@@ 3.3 Generation and Relaxation of Adsorbate-Surface Configurations
The selected surfaces were used to generate adsorbate-surface configurations using neutral fragments of *H, *OH, *OCHO and *OCH3 by means of the fairchem input generation workflow [41]. The systems were relaxed using an adsorption energy-based OCP equiformer_V2 equivariant graph neural network MLFF to obtain the relaxed geometries. The corresponding adsorption energy $E_{ads}$ follows the OCP convention $E_{ads}=E_{system}-E_{surface}-E_{adsorbate}$, where the gas phase energy $E_{adsorbate}$ is in this case the energy of radicals, and therefore hard to define. The OCP MLFFs are using the definition of Chanussot et. al. [33] , which is a linear combination of atomic energies for H, O and C, derived from H2, H2O and CO, respectively. Here, $E_{adsorbate}$ is –3.477 eV for *H, –10.552 eV for *OH, –24.917 eV for *OCH3 and –25.161 eV for *OCHO. As our work is focusing on identifying AED similarities or differences rather than specific adsorbate values, different adsorbate references would merely shift all AEDs by the same value, but will not affect our results. The force convergence criterion for the adsorbate relaxations was set to 0.03 eV/Å. The equiformer_V2 relaxations were performed using NVIDIA Ampere A100 GPUs, completing each relaxation in just a few seconds, whereas corresponding DFT calculations would require several hours on two AMD Rome 7H12 CPUs with 64 cores each. Relaxed configurations of the same material and facet were compared to identify redundancies. If two or more configurations exhibited identical adsorbate geometries (and thus energies) within a tolerance of 0.1 Å in each spatial direction, only one configuration was retained for further consideration. Nonunique configurations and their corresponding adsorption energies were excluded. With this we ensured that the remaining data represents meaningful adsorption positions and thus capture the material’s true energetic landscape.